OpenStudy (anonymous):

Write the difference in simplest form: 6/12 MINUS 1/3 =

OpenStudy (anonymous):

\[\frac{6}{12} - \frac{1}{3} \implies \frac{6}{12} - \frac{1}{3} \times \frac{4}{4} \implies \frac{6}{12} - \frac{4}{12} \implies \frac{6-4}{12} = ??\]

OpenStudy (wwe123):

mkim

OpenStudy (wwe123):

sorry 6/12=1/2 1/2+1/3 ,LCM of 2& 3 = 6 (2+3)/6 5/6

OpenStudy (anonymous):

There is minus in between and not plus..

OpenStudy (anonymous):